HYPERFINE STRUCTURE OF THE NEAR INFRARED TRANSITIONS OF CoI AND CoCl ALLY L. WONG, W. S. TAM, and \underline{A. S-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.